# Unified Framework of Forced Magnetic Reconnection and Alfvén Resonance

### v1

@@ II Model equations and boundary conditions
$\hat{g}(\hat{\phi}^{\prime\prime}-\hat{k}^{2}\hat{\phi})=i\hat{k}\hat{x}(\hat{\psi}^{\prime\prime}-\hat{k}^{2}\hat{\psi}),$ (4)
$\hat{g}\hat{\psi}=\frac{1}{S}(\hat{\psi}^{\prime\prime}-\hat{k}^{2}\hat{\psi})+i\hat{k}\hat{x}\hat{\phi},$ (5)
where prime denotes the derivative with respect to $\hat{x}$, and $\hat{g}$ and $\hat{k}$ are the Laplace and Fourier variables defined with the following conventions,
@@
for an arbitrary function $f$. Here, length scales have been normalized to the slab width $a$, and time to the Alfvén time defined by $\tau_{a}=a/v_{a}$, where $v_{a}=B_{0}/\sqrt{4\pi\rho}$ is the Alfvén speed. The Lundquist number $S=av_{a}/\eta$ is the ratio of the resistive times scale $\tau_{r}=a^{2}/\eta$ to $\tau_{a}$, $\eta$ being the magnetic diffusivity.
We look for solutions to Eqs. (4)-(5) in the domain $\hat{x}=[-1,1]$, subject to a displacement of the boundaries $\hat{\Xi}_{\pm}(\hat{k},\hat{g})$. The time-dependent solution for a given Fourier mode will then be obtained by taking the inverse Laplace transform. The form of the boundary displacement (the forcing) will be left general for much of the derivation, but it is assumed periodic in the azimuthal $y$ direction and symmetric about the midplane $x=0$. The general boundary conditions that complete our problem are therefore represented by
$\hat{\psi}(\pm 1,\hat{k},\hat{g})=\hat{\Xi}(\hat{k},\hat{g}),$ (8)
$\hat{\phi}(\pm 1,\hat{k},\hat{g})=\pm\frac{\hat{g}}{i\hat{k}}\hat{\Xi}(\hat{k},\hat{g}),$ (9)
@@ III The Time Dependent Boundary Layer Problem
Essential to the theories of both forced magnetic reconnection and Alfvén resonance is the formation of boundary layers about their resonant surfaces. We assume that boundary layers develop for $|\hat{x}|\ll 1$. This means that the regime we are exploring for the Alfvén resonance is that of a sub-Alfvénic resonant frequency $\omega_{0}\tau_{a}\ll 1$. Therefore, for $S\gg 1$, the plasma can be assumed to be governed by ideal MHD (or $S=\infty$) except in the immediate vicinity of the boundary layer, which contains large gradients. In the boundary layer, or inner layer, resistive MHD must be considered, however, because of the large gradients, one can assume that $\partial^{2}/\partial\hat{x}^{2}\gg\hat{k}^{2}$. In this way, equations (4) and (5) can be solved by breaking the domain into two regions: an ideal “outer region” away from the boundary layer, and an “inner layer” which contains all the physics of the problem but is treated as one dimensional. These regions are then asymptotically matched for a fully general solution valid in the entire domain.
@@ III.1 The Outer Region
$\hat{\psi}^{\prime\prime}=\hat{k}^{2}\hat{\psi},$ (10)
$\hat{\phi}=\frac{\hat{g}}{i\hat{k}\hat{x}}\hat{\psi}.$ (11)
The general solution to Eq. (10) that satisfies the boundary condition given by Eq. (8) and is even in $\hat{x}$ is given by [2]
$\hat{\psi}^{out}=\hat{\psi}_{0}(\hat{k},\hat{g})\left[\cosh(\hat{k}\hat{x})-\frac{\sinh(\hat{k}|\hat{x}|)}{\tanh(\hat{k})}\right]+\hat{\Xi}(\hat{k},\hat{g})\frac{\sinh(\hat{k}|\hat{x}|)}{\sinh(\hat{k})},$ (12)
where $\hat{\psi}_{0}(\hat{k},\hat{g})=\hat{\psi}^{out}(\hat{x}=0)$. The outer solution for $\hat{\phi}$ follows directly from Eq. (11). The solution (12) consists of two contributions: a tearing mode contribution satisfying homogeneous boundary conditions that takes a value of amplitude $\hat{\psi}_{0}(\hat{k},\hat{g})$ at the rational surface $\hat{x}=0$, representing the reconnected flux, and an ideal contribution describing a “screened” plasma response that vanishes at $\hat{x}=0$ given by $\hat{\Xi}$ at the boundary. As can be seen from Eq. (12), $\hat{\psi}^{out}$ exhibits a jump in its derivative which we define here as
$\Delta\Psi^{\prime}\equiv-\frac{2\hat{k}}{\tanh(\hat{k})}\hat{\psi}_{0}+\frac{2\hat{k}}{\sinh(\hat{k})}\hat{\Xi}.$ (13)
@@
$\hat{\psi}^{out}(\hat{x}\rightarrow 0)=\hat{\psi}_{0}+\frac{1}{2}\Delta\Psi^{\prime}|\hat{x}|,$ (14)
$\hat{\phi}^{out}(\hat{x}\rightarrow 0)=\frac{\hat{g}}{i\hat{k}}\hat{\psi}_{0}\frac{1}{\hat{x}}\left(1+\frac{1}{2}\frac{\Delta\Psi^{\prime}}{\hat{\psi}_{0}}|\hat{x}|\right).$ (15)
The limit of the outer solution for $\hat{x}\rightarrow 0$ expressed by Eq. (14) and (15), provides the boundary condition for the inner layer solution as it tends away from the resonance. We anticipate that the inner layer equations can be solved for $\hat{\phi}$ by applying a Fourier transform with respect to the variable $\hat{x}$ [19, 20]. With that in mind, it follows from Eq. (15) that the asymptotic behavior of the Fourier transform of ${\phi}^{out}(\hat{x})$, denoted as $\tilde{\phi}^{out}(\hat{\theta})$, at large $\hat{\theta}$ is given by
$\tilde{\phi}^{out}(\hat{\theta}\rightarrow\infty)=-\frac{\hat{g}}{\hat{k}}\frac{\hat{\psi_{0}}}{2}\frac{1}{\hat{\theta}}\left(|\hat{\theta}|+\frac{1}{\pi}\Delta_{g}\right),$ (16)
where from now on we use the tilde $\tilde{[\ ]}$ to denote functions in $\hat{\theta}$ space, and
@@ III.2 The Inner Layer
$\hat{g}\hat{\phi}^{\prime\prime}=i\hat{k}\hat{x}\hat{\psi}^{\prime\prime},$ (18)
$\hat{\psi}=\frac{1}{\hat{g}S}\hat{\psi}^{\prime\prime}+\frac{i\hat{k}\hat{x}}{\hat{g}}\hat{\phi},$ (19)
respectively. Just like equations (4)-(5), the system of equations (18)-(19) admit solutions for $\hat{\psi}$ and $\hat{\phi}$ with a defined parity. In the following we seek solutions with $\hat{\psi}$ even in $\hat{x}$ and $\hat{\phi}$ odd, and we will show only the $\hat{x}\geq 0$ part of the solutions. The solution in the entire domain can then be obtained by extension, according to the appropriate parity.
To make the above set of equations analytically tractable, we Fourier transform in the $\hat{x}$ direction [21] obtaining the following (lower order) differential equations in terms of $\hat{\theta}$,
$\hat{\theta}^{2}\tilde{\phi}=-\frac{\hat{k}}{\hat{g}}\frac{\partial}{\partial\hat{\theta}}(\hat{\theta}^{2}\tilde{\psi}),$ (20)
$\left(1+\frac{\hat{\theta}^{2}}{\hat{g}S}\right)\tilde{\psi}=-\frac{\hat{k}}{\hat{g}}\frac{\partial}{\partial\hat{\theta}}\tilde{\phi}.$ (21)
From the definition $J_{z}=-\nabla^{2}\psi\approx\hat{\theta}^{2}\tilde{\psi}$ and equation (21), we derive the following expression for $\tilde{J}_{z}$ in terms of $\tilde{\phi}$,
$\tilde{J}_{z}=-\frac{(S\hat{k})^{2/3}}{Q}\frac{\bar{\theta}^{2}}{(1+\frac{\bar{\theta}^{2}}{Q})}\frac{d}{d\bar{\theta}}\tilde{\phi},$ (22)
where we have introduced the normalized variables
$Q=\frac{\hat{g}S^{1/3}}{\hat{k}^{2/3}},\quad\bar{\theta}=\frac{\hat{\theta}}{(S\hat{k})^{1/3}}.$ (23)
Next, by dividing equation (20) by $\bar{\theta}^{2}$, differentiating it with respect to $\bar{\theta}$, and making use of Eq. (22), we find an expression entirely in terms of the current density
$\bar{\theta}^{2}\frac{d}{d\bar{\theta}}\left(\frac{1}{\bar{\theta}^{2}}\frac{d}{d\bar{\theta}}\tilde{J_{z}}\right)=Q^{2}\left(1+\frac{\bar{\theta}^{2}}{Q}\right)\tilde{J_{z}}.$ (24)
@@
$z=\sqrt{Q}\bar{\theta}^{2}=\frac{\sqrt{Q}\hat{\theta}^{2}}{(S\hat{k})^{2/3}},$ (25)
and with the ansatz $\tilde{J}_{z}(z)=e^{-\frac{z}{2}}f(z)$ (based on the expected behavior of the solution by taking the limit $z\gg 1$ of Eq. (24)), Eq. (24) takes the form of Kummer’s confluent hypergeometric equation for the function $f(z)$:
$zf^{\prime\prime}+\left(-\frac{1}{2}-z\right)f^{\prime}-\frac{(Q^{3/2}-1)}{4}f=0.$ (26)
@@
where $A$ is an amplitude to be determined via asymptotic matching. By using the recurrence properties of the exponential and of the function $U(a,b,z)$, together with equation (20) upon substitution of $\hat{J}_{z}=\hat{\theta}^{2}\tilde{\psi}$, we determine the inner layer solution for $\tilde{\phi}$ as:
$\tilde{\phi}^{in}=\frac{2A}{(S\hat{k})^{2/3}Q^{1/4}}\frac{e^{-\frac{z}{2}}}{\sqrt{z}}\left[U\left(\frac{(Q^{3/2}-1)}{4},\frac{1}{2},z\right)-\frac{1}{2}U\left(\frac{(Q^{3/2}-1)}{4},-\frac{1}{2},z\right)\right].$ (28)
In the limit of small $\hat{\theta}$ (or large $\hat{x}$), the solution to be asymptotically matched with Eq. (16) is
$\tilde{\phi}^{in}_{\hat{\theta}\rightarrow 0}=\frac{2\sqrt{\pi}A}{(S\hat{k})^{2/3}Q^{1/4}}\frac{1}{\hat{\theta}}\left(\frac{Q^{5/4}}{(Q^{3/2}+1)}\frac{(S\hat{k})^{1/3}}{\Gamma\left(\frac{Q^{3/2}+1}{4}\right)}-\frac{2}{\Gamma\left(\frac{Q^{3/2}-1}{4}\right)}\hat{\theta}\right)$ (29)
@@ IV The Matched Inner Layer Solution
$A=\frac{\hat{\psi}_{0}(\hat{k},\hat{g})}{8\sqrt{\pi}}Q^{5/4}(S\hat{k})^{1/3}\Gamma\left(\frac{(Q^{3/2}-1)}{4}\right),$ (30)
@@
$\hat{\psi}_{0}(\hat{k},\hat{g})=\hat{\Xi}(\hat{k},\hat{g})\frac{\frac{2\hat{k}}{\sinh[\hat{k}]}}{\Delta_{g}+\frac{2\hat{k}}{\tanh[\hat{k}]}}\equiv\hat{\Xi}(\hat{k},\hat{g})\frac{\Delta_{f}}{\Delta_{g}-\Delta_{0}},$ (31)
@@
$\Delta_{g}=-2\pi(S\hat{k})^{1/3}\frac{Q^{5/4}}{(Q^{3}-1)}\frac{\Gamma\left(\frac{(Q^{3/2}-1)}{4}+1\right)}{\Gamma\left(\frac{(Q^{3/2}-1)}{4}+\frac{1}{2}\right)}.$ (32)
Here $\Delta_{f}=\frac{2\hat{k}}{\sinh{\hat{k}}}$ and $\Delta_{0}=-\frac{2\hat{k}}{\tanh{\hat{k}}}$, with $\Delta_{0}$ being the conventional jump of the derivative of $\hat{\psi}^{out}$ at the origin (the $\Delta^{\prime}$ parameter of the tearing mode).
Equation (28), together with Eqs. (30)-(32), completely describes the stream function for both forced magnetic reconnection and Alfvén resonance, and it includes the full time evolution from the early stages, $\hat{t}\ll S^{1/3}$, when the evolution is essentially ideal and the boundary layer is being formed, all the way to the long-time behavior determined by resistivity.
We note that Eq. (28) includes as special cases the eigenfunctions that describe the ideal kink, the resistive kink (also referred to as non constant-psi regime for reconnection in slab geometry) and the ordinary tearing mode (constant-psi regimes) under appropriate asymptotic expansions [23, 21, 24, 20] [^1] We note a typo on equation (31) of [20].. The important difference is that here we are solving for the time-dependent problem rather than the eigenvalue problem and, therefore, $\hat{g}$ is a variable; additionally, we include the effect of a general forcing function at the boundary through the function $\hat{\Xi}$ by allowing not only (forced) reconnection, but also Alfvén resonances.
The solution given in Eq. (28) will now be considered for two distinct stages of the dynamical evolution of the system. The early-time ideal regime corresponds to the limit $Q\gg 1$ ($\hat{t}\ll S^{1/3}$). The time-asymptotic resistive regime, when resistivity has allowed for the formation of a finite width boundary layer, corresponds to the limit $Q\ll 1$ ($\hat{t}\gg S^{1/3}$). These two stages, which forced reconnection and Alfvén resonance have in common, are naturally recovered through two asymptotic expansions of the general solution (28) with respect to the parameter $Q$. The transition between the ideal and the resistive regimes occurs for $Q\sim 1$, or for times $\hat{t}\sim S^{1/3}$. Thus, with Eq. (28) we recover the expected time scales.
@@
$\hat{\Xi}(\hat{k},\hat{g})=\frac{\Xi_{0}}{2}\frac{\hat{g}}{(\hat{g}^{2}+\hat{\omega}_{0}^{2})}\left[\delta(\hat{k}-\hat{k}_{0})+\delta(\hat{k}+\hat{k}_{0})\right].$ (34)

### v2

@@ II Model equations and boundary conditions
$g(\hat{\phi}^{\prime\prime}-k^{2}\hat{\phi})=ikx(\hat{\psi}^{\prime\prime}-k^{2}\hat{\psi}),$ (4)
$g\hat{\psi}=\frac{1}{S}(\hat{\psi}^{\prime\prime}-k^{2}\hat{\psi})+ikx\hat{\phi},$ (5)
where prime denotes the derivative with respect to $x$, and $g$ and $k$ are the Laplace and Fourier variables defined with the following conventions,
@@
for an arbitrary function $f$. For ease of notation, hereafter we will denote with hats Laplace transforms (therefore functions of $g$), which may or may not depend on the Fourier variable $k$. In case of ambiguity, we will indicate explicitly the variables upon which a given function depends. Here, length scales have been normalized to the slab width $a$, and time to the Alfvén time defined by $\tau_{a}=a/v_{a}$, where $v_{a}=B_{0}/\sqrt{4\pi\rho}$ is the Alfvén speed. The Lundquist number $S=av_{a}/\eta$ is the ratio of the resistive times scale $\tau_{r}=a^{2}/\eta$ to $\tau_{a}$, $\eta$ being the magnetic diffusivity.
We look for solutions to Eqs. (4)-(5) in the domain $x=[-1,1]$, subject to a displacement of the boundaries $\hat{\Xi}_{\pm}(k,g)$. The time-dependent solution for a given Fourier mode will then be obtained by taking the inverse Laplace transform. The form of the boundary displacement (the forcing) will be left general for much of the derivation, but it is assumed periodic in the azimuthal $y$ direction and symmetric about the midplane $x=0$. The general boundary conditions that complete our problem are therefore represented by
$\hat{\psi}(\pm 1,k,g)=\hat{\Xi}(k,g),$ (8)
$\hat{\phi}(\pm 1,k,g)=\pm\frac{g}{ik}\hat{\Xi}(k,g),$ (9)
@@ III The Time Dependent Boundary Layer Problem
Essential to the theories of both forced magnetic reconnection and Alfvén resonance is the formation of boundary layers about their resonant surfaces. We assume that boundary layers develop for $|x|\ll 1$. This means that the regime we are exploring for the Alfvén resonance is that of a sub-Alfvénic resonant frequency $\omega_{0}\tau_{a}\ll 1$. Therefore, for $S\gg 1$, the plasma can be assumed to be governed by ideal MHD (or $S=\infty$) except in the immediate vicinity of the boundary layer, which contains large gradients. In the boundary layer, or inner layer, resistive MHD must be considered, however, because of the large gradients, one can assume that $\partial^{2}/\partial x^{2}\gg k^{2}$. In this way, equations (4) and (5) can be solved by breaking the domain into two regions: an ideal “outer region” away from the boundary layer, and an “inner layer” which contains all the physics of the problem but is treated as one dimensional. These regions are then asymptotically matched for a fully general solution valid in the entire domain.
@@ III.1 The Outer Region
$\hat{\psi}^{\prime\prime}=k^{2}\hat{\psi},$ (10)
$\hat{\phi}=\frac{g}{ikx}\hat{\psi}.$ (11)
The general solution to Eq. (10) that satisfies the boundary condition given by Eq. (8) and is even in $x$ is given by [2]
$\hat{\psi}^{out}=\hat{\psi}_{0}(k,g)\left[\cosh(kx)-\frac{\sinh(k|x|)}{\tanh(k)}\right]+\hat{\Xi}(k,g)\frac{\sinh(k|x|)}{\sinh(k)},$ (12)
where $\hat{\psi}_{0}(k,g)=\hat{\psi}^{out}(x=0)$. The outer solution for $\hat{\phi}$ follows directly from Eq. (11). The solution (12) consists of two contributions: a contribution satisfying homogeneous boundary conditions (no forcing) that takes a time dependent value of amplitude $\hat{\psi}_{0}(k,g)$ at the rational surface $x=0$ and an ideal contribution describing a “screened” plasma response that vanishes at $x=0$ that satisfies the boundary condition. As can be seen from Eq. (12), $\hat{\psi}^{out}$ exhibits a jump in its derivative which we define here as
$\Delta\Psi^{\prime}\equiv-\frac{2k}{\tanh(k)}\hat{\psi}_{0}+\frac{2k}{\sinh(k)}\hat{\Xi}.$ (13)
@@
$\hat{\psi}^{out}(x\rightarrow 0)=\hat{\psi}_{0}+\frac{1}{2}\Delta\Psi^{\prime}|x|,$ (14)
$\hat{\phi}^{out}(x\rightarrow 0)=\frac{g}{ikx}\hat{\psi}_{0}\left(1+\frac{1}{2}\frac{\Delta\Psi^{\prime}}{\hat{\psi}_{0}}|x|\right).$ (15)
The limit of the outer solution for $x\rightarrow 0$ expressed by Eq. (14) and (15), provides the boundary condition for the inner layer solution as it tends away from the resonance. We anticipate that the inner layer equations can be solved for $\hat{\phi}$ by applying a Fourier transform with respect to the variable $x$ [19, 20]. With that in mind, it follows from Eq. (15) that the asymptotic behavior of the Fourier transform of ${\phi}^{out}(x)$, denoted as $\tilde{\phi}^{out}(\theta)$, at large $\theta$ is given by
$\tilde{\phi}^{out}(\theta\rightarrow\infty)=-\frac{g}{k}\frac{\hat{\psi_{0}}}{2}\frac{1}{\theta}\left(|\theta|+\frac{1}{\pi}\Delta_{g}\right),$ (16)
where from now on we use the tilde $\tilde{[\ ]}$ to denote functions in $\theta$ space, and
@@ III.2 The Inner Layer
$g\hat{\phi}^{\prime\prime}=ikx\hat{\psi}^{\prime\prime},$ (18)
$\hat{\psi}=\frac{1}{gS}\hat{\psi}^{\prime\prime}+\frac{ikx}{g}\hat{\phi},$ (19)
respectively. Just like equations (4)-(5), the system of equations (18)-(19) admit solutions for $\hat{\psi}$ and $\hat{\phi}$ with a defined parity. In the following we seek solutions with $\hat{\psi}$ even in $x$ and $\hat{\phi}$ odd, and we will show only the $x\geq 0$ part of the solutions. The solution in the entire domain can then be obtained by extension, according to the appropriate parity.
To make the above set of equations analytically tractable, we Fourier transform in the $x$ direction. This approach allows one to reduce the original fourth order system of differential equations with respect to the variable $x$, to a second order system in $\theta$ that is analytically solvable [21]. In this way, the differential equations in terms of $\theta$ are
$\theta^{2}\tilde{\phi}=-\frac{k}{g}\frac{\partial}{\partial\theta}(\theta^{2}\tilde{\psi}),$ (20)
$\left(1+\frac{\theta^{2}}{gS}\right)\tilde{\psi}=-\frac{k}{g}\frac{\partial}{\partial\theta}\tilde{\phi}.$ (21)
From the definition $J_{z}=-\nabla^{2}\psi\approx\theta^{2}\tilde{\psi}$ and equation (21), we derive the following expression for $\tilde{J}_{z}$ in terms of $\tilde{\phi}$,
$\tilde{J}_{z}=-\frac{(Sk)^{1/3}}{Q}\frac{\theta^{2}}{(1+\frac{\theta^{2}}{(Sk)^{2/3}Q})}\frac{d}{d\theta}\tilde{\phi},$ (22)
where we have introduced the normalized Laplace variable
$Q=\frac{gS^{1/3}}{k^{2/3}}.$ (23)
Next, by dividing equation (20) by $\theta^{2}$, differentiating it with respect to $\theta$, and making use of Eq. (22), we find an expression entirely in terms of the current density
$\theta^{2}\frac{d}{d\theta}\left(\frac{(Sk)^{2/3}}{\theta^{2}}\frac{d}{d\theta}\tilde{J_{z}}\right)=Q^{2}\left(1+\frac{\theta^{2}}{(Sk)^{2/3}Q}\right)\tilde{J_{z}}.$ (24)
@@
$z=\frac{\sqrt{Q}\theta^{2}}{(Sk)^{2/3}},$ (25)
and with the ansatz $\tilde{J}_{z}(z)=e^{-\frac{z}{2}}\tilde{f}(z)$ (based on the expected behavior of the solution by taking the limit $z\gg 1$ of Eq. (24)), Eq. (24) takes the form of Kummer’s confluent hypergeometric equation for the function $\tilde{f}(z)$:
$z\tilde{f}^{\prime\prime}+\left(-\frac{1}{2}-z\right)\tilde{f}^{\prime}-\frac{(Q^{3/2}-1)}{4}\tilde{f}=0.$ (26)
@@
where $A$ is an amplitude to be determined via asymptotic matching. By using the recurrence properties of the exponential and of the function $U(a,b,z)$, together with equation (20) upon substitution of $\tilde{J}_{z}=\theta^{2}\tilde{\psi}$, we determine the inner layer solution for $\tilde{\phi}$ as:
$\tilde{\phi}^{in}=\frac{2A}{(Sk)^{2/3}Q^{1/4}}\frac{e^{-\frac{z}{2}}}{\sqrt{z}}\left[U\left(\frac{(Q^{3/2}-1)}{4},\frac{1}{2},z\right)-\frac{1}{2}U\left(\frac{(Q^{3/2}-1)}{4},-\frac{1}{2},z\right)\right].$ (28)
In the limit of small $\theta$ (or large $x$), the solution to be asymptotically matched with Eq. (16) is
$\tilde{\phi}^{in}_{\theta\rightarrow 0}=\frac{2\sqrt{\pi}A}{(Sk)^{2/3}Q^{1/4}}\frac{1}{\theta}\left(\frac{Q^{5/4}}{(Q^{3/2}+1)}\frac{(Sk)^{1/3}}{\Gamma\left(\frac{Q^{3/2}+1}{4}\right)}-\frac{2}{\Gamma\left(\frac{Q^{3/2}-1}{4}\right)}\theta\right)$ (29)
@@ IV The Matched Inner Layer Solution
$A=\frac{\hat{\psi}_{0}(k,g)}{8\sqrt{\pi}}Q^{5/4}(Sk)^{1/3}\Gamma\left(\frac{(Q^{3/2}-1)}{4}\right),$ (30)
@@
$\hat{\psi}_{0}(k,g)=\hat{\Xi}(k,g)\frac{\frac{2k}{\sinh[k]}}{\Delta_{g}+\frac{2k}{\tanh[k]}}\equiv\hat{\Xi}(k,g)\frac{\Delta_{f}}{\Delta_{g}-\Delta_{0}},$ (31)
@@
$\Delta_{g}=-2\pi(Sk)^{1/3}\frac{Q^{5/4}}{(Q^{3}-1)}\frac{\Gamma\left(\frac{(Q^{3/2}-1)}{4}+1\right)}{\Gamma\left(\frac{(Q^{3/2}-1)}{4}+\frac{1}{2}\right)}.$ (32)
Here $\Delta_{f}=\frac{2k}{\sinh{k}}$ and $\Delta_{0}=-\frac{2k}{\tanh{k}}$, with $\Delta_{0}$ being the conventional jump of the derivative of $\hat{\psi}^{out}$ at the origin (the $\Delta^{\prime}$ parameter of the tearing mode).
Equation (28), together with Eqs. (30)-(32), completely describes the stream function for both forced magnetic reconnection and Alfvén resonance, and it includes the full time evolution from the early stages, $t\ll S^{1/3}$, when the evolution is essentially ideal and the boundary layer is being formed, all the way to the long-time behavior determined by resistivity.
We note that Eq. (28) includes as special cases the eigenfunctions that describe the ideal kink, the resistive kink (also referred to as non constant-psi regime for reconnection in slab geometry) and the ordinary tearing mode (constant-psi regimes) under appropriate asymptotic expansions [23, 21, 24, 20] [^1] We note a typo on equation (31) of [20].. The important difference is that here we are solving for the time-dependent problem rather than the eigenvalue problem and, therefore, $g$ is a variable; additionally, we include the effect of a general forcing function at the boundary through the function $\hat{\Xi}$ by allowing not only (forced) reconnection, but also Alfvén resonances.
The solution given in Eq. (28) will now be considered for two distinct stages of the dynamical evolution of the system. The early-time ideal regime corresponds to the limit $Q\gg 1$ ($t\ll S^{1/3}$). The time-asymptotic resistive regime, when resistivity has allowed for the formation of a finite width boundary layer, corresponds to the limit $Q\ll 1$ ($t\gg S^{1/3}$). These two stages, which forced reconnection and Alfvén resonance have in common, are naturally recovered through two asymptotic expansions of the general solution (28) with respect to the parameter $Q$. The transition between the ideal and the resistive regimes occurs for $Q\sim 1$, or for times $t\sim S^{1/3}$. Thus, with Eq. (28) we recover the expected time scales.
@@
$\hat{\Xi}(k,g)=\frac{\Xi_{0}}{2}\frac{g}{(g^{2}+\omega_{0}^{2})}\left[\delta(k-k_{0})+\delta(k+k_{0})\right].$ (34)